Solve for X:
x^2+4x+4=7
The answer I recieved was:
X=.645
My teacher has said this is wrong and is not in algebra format. Can someone please help me with the answer and the correct format? Thank you so much for the wonderful help this site brings me and many others.

Found 3 solutions by checkley71, stanbon, Edwin McCravy:
Answer by checkley71(8403)

x^2+4x+4=7
x^2+4x+4-7=0
x^2+4x-3=0 using the quadratic equation we get
x=(-b+-sqrt[b^2-4ac])/2a
x=(-4+-sqrt4^2-4*1*-3])/2*1
x=(-4+-sqrt16+12])/2
x=(-4+-sqrt28)/2
x=(-4+-5.29)/2
x=(-4+5.29)/2
x=-1.29/2
x=.645 solution
x=(-4-5.29)/2
x=-9.29/2
x=-4.645 solution
----------------------------------------------------
HAPPY NEW YEAR

Answer by stanbon(75887)

Solve for X:
x^2+4x+4=7
x^2+4x-3=0
x=[-4+-sqrt(16-4*-3]/2
x=[-2+-2sqrt7]/2
x=[-1+-sqrt7]
x=-1+sqrt7] or x=[-1-sqrt7]
Cheers,
Stan H.

Answer by Edwin McCravy(20055)

Solve for X:
x^2+4x+4=7
The answer I recieved was:
X=.645
My teacher has said this is wrong and is not in
algebra format. Can someone please help me with 
the answer and the correct format? Thank you so 
much for the wonderful help this site brings me 
and many others.

Get 0 on the right by adding -7 to both sides

           x² + 4x + 4 = 7
                    -7  -7
           ---------------
           x² + 4x - 3 = 0

That won't factor, so we use the quadratic formula:
                  ______ 
            -b ± Öb²-4ac
        x = —————————————
                2a 

where a = 1; b = 4; c = -3

                     _____________ 
             -(4) ± Ö(4)²-4(1)(-3)
        x = ————————————————————————
                     2(1) 
                   _____ 
             -4 ± Ö16+12
        x = —————————————
                 2

                   __ 
             -4 ± Ö28
        x = ——————————
                2

                   ___ 
             -4 ± Ö4·7
        x = ———————————
                2 

                   _
            -4 ± 2Ö7
        x = —————————
                2

                     _
            -4     2Ö7
        x = ——— ± —————
             2      2
                  _
        x = -2 ± Ö7 
                       _
Using the +, x = -2 + Ö7, which
is one solution and equals about .6457513111,
so your answer of .645 is close to that, but rounded 
off to thousandths should be .646, not .645.
                       _ 
Using the -, x = -2 - Ö7, which
is the other solution and equals about -4.645751311

However the solutions that your teacher wants are
                  _
        x = -2 ± Ö7

Edwin